OpenStudy (anonymous):

If beaker A contains a 15% solution of Lemonade and beaker B contains a 20% solution of Lemonade, how much from each beaker shoud be usede to make 10 liters of a 16% solution of lemonade

OpenStudy (nikato):

So we know there are 10 liters of the mix solution which is 16%= 10(0.16) Lets say beaker A has x liters and is 15%= 0.15x And beaker B has 10-x liters and is 20% = (10-x)0.20 Beaker A + Beaker B= mix So 0.15x + (10-x)0.20 = 10(0.16) Now try to solve for x

OpenStudy (anonymous):

ok I will

OpenStudy (dumbcow):

you can get rid of decimals if you multiply everything by 100

OpenStudy (anonymous):

I got 20 for x

OpenStudy (nikato):

No.

OpenStudy (nikato):

0.15x + (10-x)0.20 = 10(0.16) 0.15x + 2 -0.2x= 1.6 -2. -2 -------------------------- 0.15x -0.2x = -0.4 -0.05x = -0.4 ------ ---- -0.05. -0.05 x=?

OpenStudy (anonymous):

I don't know why but for some reason I thought 10 time .2 was 20 sorry

OpenStudy (nikato):

It's ok. Everyone makes mistakes